OpenStudy (kirbykirby):

I have this change of indices: \[\sum_{j=0}^{\infty} \sum_{k=0}^{\infty} \frac{z^j}{j!} \frac{w^k}{k!} = \sum_{n=0}^{\infty} \sum_{k=0}^{n} \frac{z^{n-k}}{(n-k)!}\]\frac{w^k}{k!} I'm not too sure how to get the k=0 to n inner indices.

OpenStudy (kirbykirby):

Oops \[\sum_{j=0}^{\infty} \sum_{k=0}^{\infty} \frac{z^j}{j!} \frac{w^k}{k!} = \sum_{n=0}^{\infty} \sum_{k=0}^{n} \frac{z^{n-k}}{(n-k)!} \frac{w^k}{k!}\]

jigglypuff314 (jigglypuff314):

Hello! This is the feedback group of this site, you would get better help if you posted this in the Mathematics group http://openstudy.com/study#/groups/Mathematics Good Luck! :)

OpenStudy (xmoses1):

You could ask this in the \( \LaTeX~Practicing \) section. I bet someone there will know

OpenStudy (xmoses1):

Oh wait never mind listen to @jigglypuff314 lol

OpenStudy (kirbykirby):

Oh I honestly don't know how I got to this part... I swear I was on the Math section :O

OpenStudy (xmoses1):

XD it happens lol dont sweat it!